# What Makes LLM Agent Simulations Useful for Policy? Insights From an Iterative Design Engagement in Emergency Preparedness

### v1

@@ 5.6. Fifth Iteration: Exploring Better Policy Implementation
5.6.1. Simulation Goals and Design
For this iteration, we used the validated system to explore simulation outputs for novel scenarios and policy implementation improvements identified by policymakers. Building directly on their feedback, we implemented a set of scenario variations that combined:
  • two emergency types (severe weather and bomb threat)
  • evacuation announcements with and without specific location information,
  • alternative coordinator positioning strategies
  • the addition of an extra exit
We tested severe weather scenarios where all exits remained accessible and bomb threat scenarios where specific areas required targeted evacuations away from threat locations. Announcement variations included location-specific messages informing agents about threat locations versus general evacuation instructions. Coordinator layouts compared existing positioning with alternative distributions such as placing coordinators closer to seating sections rather than concentrated near exits.
These variations were designed to reflect the specific “what-if” questions raised by the policymaker team and to test the robustness of different preparedness strategies. To evaluate the impact of these policy implementation options, we measured evacuation efficiency by the time required for 80% of agents to complete their evacuation—the outcome metric that policymakers consistently emphasized as most relevant for assessing operational success.
5.6.2. Policy Implementation
This iteration in August 2025 produced three concrete policy implementation proposals that emerged directly from simulation findings:
@@ 5.6.2. Policy Implementation
    Opening a previously unused exit at the northwestern corner of the stadium to reduce evacuation bottlenecks
@@
    Training coordinators to guide people located both near to and far away from localized threats rather than focusing only on nearby individuals
@@
    Developing differentiated evacuation protocols for various emergency types and incorporating these distinctions into coordinator training materials.
The policymaker team adopted Proposals 2 and 3, incorporating the expanded coordinator guidance protocols and differentiated emergency procedures into their standard operating protocols. However, they also noted that in man-made threat scenarios—such as bomb or active-shooter incidents—operational authority shifts to local police. In such cases, the preparedness team’s role would be limited, constraining their ability to fully implement these process changes. For Proposal 1, policymakers initiated a formal feasibility assessment process to evaluate the structural and logistical requirements for opening the additional exit, representing the standard institutional pathway for implementing infrastructure changes.
The after-action report in August 2025 documented these outcomes, emphasizing both sustained institutional integration of the simulations and their broader significance for large-scale public safety planning. Policymakers explicitly committed to ongoing development of the system as a long-term resource for preparedness and training.
6. Findings
In our study, “usefulness” did not follow a linear progression from technical sophistication to institutional adaptation. Rather, it emerged through an iterative, stakeholder-engaged design process in which policymakers gradually built trust in simulation outputs, learned the boundaries of what LLM agent simulations can and cannot do, and re-calibrated their expectations and criteria based on this hard-won trust and understanding.
Analyzing this process, we identified five key insights on developing LLM agent simulations that are “useful” for policy by supporting implementation:
  1. (1)
    Validation Filter: Usefulness requires scenarios that can be validated
  2. (2)
    Trust Bootstrap: Trust from validating mundane scenarios extends to novel ones
  3. (3)
    “Fix-It” Response: “Wrong” simulations surface tacit domain knowledge from policymakers
@@ 6. Findings
We discuss each of these insights in the subsections that follow.
@@ 6.1. Validation Filter: Usefulness Requires Scenarios That Can be Validated
In our study, LLM agent simulations were not deemed useful in general, but only in scenarios where outputs could be validated against policymakers’ prior experience or ground-truth data. Policymakers did not state this requirement explicitly at the outset, but over time it became clear that observable, verifiable scenarios were essential entry points for building confidence in simulation systems. Only after this baseline validation did they begin to engage with outputs for less well-understood or hypothetical emergency scenarios.
We describe this as a “validation filter”: LLM agent simulations can be perceived as useful only in application domains where results from well-understood scenarios can be checked by multiple stakeholders together. This filter is particularly important given the stochastic nature of LLM outputs: both policymakers and developers needed to see where the system aligned with known realities before exploring how it might generate value in untested situations.
This observation emerged immediately in our first iteration, which simulated social media misinformation dynamics. Despite acknowledging misinformation as a pressing concern, policymakers expressed fundamental skepticism about validation possibilities. As P3 explained:
It’s very hard for us to access social media data in previous emergencies, and some important communication may happen at private Discord channels… And then you have people in other countries who may be having a conversation that’s blowing up way beyond what is actually happening here… we don’t know, I mean, I don’t know how to communicate directly with some of our audiences overseas.
@@ 6.2. Trust Bootstrap: Trust from Validating Mundane Scenarios Extends to Novel Ones
The validation filter creates a paradox: Simulations are most valuable for situations that are difficult to observe or anticipate (Park et al., 2022, 2023; Binz et al., 2025), yet they can only be trusted if grounded in scenarios that are already well understood. We found that this paradox could be addressed through a “trust bootstrap” process.
In our study, once trust was established through successful validation in mundane scenarios (in our case, attendees exiting the stadium after commencement), that trust could be extended—or “bootstrapped”—to more novel and speculative scenarios where robust validation was harder or practically impossible. This process has become especially salient with recent advances in LLMs, which allow simulation systems to adapt to natural-language scenarios while building on validated setups.
This dynamics was reflected in P1’s comment during the fourth iteration:
…when you showed a baseline evacuation without anything major, it mirrored the 20 minutes of what you all experienced with us with just the regular evacuation. And your simulation mirrored what we experience every time we have it. So that gives us some baseline trust, if you will, to see that it is connecting to what experiences we already had.
@@
We also found that iterative process was essential for developing trust because it allowed policymakers to respond to simulations with “fix-it” reactions. Even when policymakers were deeply familiar with their domain, they did not initially know which details were critical for interpreting a simulation until they saw something that looked off. Incomplete, simplistic, or otherwise “wrong” simulations therefore served as productive technology probes (Hutchinson et al., 2003a). By reacting to outputs that seemed unrealistic or missing, policymakers revealed hidden assumptions and unspoken expertise, which in turn guided refinements that brought the simulations closer to institutional relevance.
@@
For example, our original commencement evacuation simulation did not consider people’s accessibility situations. When observing that LLM agents move in identical speed smoothly, P5 immediately flagged this omission:
@@
Did you think about like people with disabilities and they might want to exit like through the one that has like the ramp or anything like that? Did you guys think about those factors?
@@
This requirement—accounting for accessibility situations—did not emerge until after policymakers saw a simulation that failed to include it. As P1 later reflected:
@@
In this way, unrealistic or incomplete outputs did not simply highlight errors—they revealed hidden assumptions about crowd psychology and institutional protocols. These insights were not captured through direct questioning but surfaced only when policymakers engaged with “wrong” simulations.
LLM agent simulations consist of two core elements: the agents themselves and the environments in which they interact (Sumers et al., 2023). In our early iterations, we focused primarily on designing agents and leveraging LLM capabilities. However, policymakers consistently pointed out that without sufficient nuance in the interaction environment, the simulations lacked credibility.
We were thinking that this evacuation is due to severe weather, which would be a fairly benign reason. But it could also be because of a bomb threat, and that opens up a whole different conversation.
In academic demonstrations, developers often concentrate on designing LLM agents. However, our findings show that the interaction environment matters just as much. Context reinforced utility by anchoring simulations in recognizable realities, and it sparked imagination about new applications. Missing or misaligned contexts could cause temporary dips in confidence, but these gaps also triggered “fix-it” responses that revealed additional requirements (see Sec. 6.3).
Our iterative process ultimately revealed a coevolution between policy requirements and simulation capabilities. While we initially aimed to design simulations that could directly inform “policy,” policymakers clarified that our assumption was misplaced. P1 stated:
These are process changes, not policy or concept of policy is different. Our policy is to implement safe procedures to make sure that we protect life, property. So, policy has a context of what we intend to do… So, our policy is the same. What we’re trying to do is to find ways to fulfill the policy, which is protecting life.
This distinction—between stable, high-level policy commitments and the evolving processes used to implement them—proved crucial for understanding where simulations could have impact. Policy represents overarching commitments that remain constant across scenarios, while policy implementation, such as processes and procedures, changes in response to context.
Usefulness therefore emerges recursively. Simulations are not one-directional tools for generating new policies; rather, they become valuable when policymakers use them to refine how existing policy goals can be operationalized in practice. In this way, both problems and solutions evolve iteratively alongside simulation development, creating a feedback loop between simulation design and policy implementation.
@@ 7.1. Are LLM Agent Simulations Useful for Policy?
Scholars remain divided on the usefulness of LLM agent simulations. This debate reflects a fundamental tension captured in Bill Buxton’s distinction between “building the thing right” versus “building the right thing (Buxton, 2010).” Proponents emphasize “building the thing right,” focusing on technical sophistication and agent architectures (Park et al., 2023, 2024; Li et al., 2024a; Horton, 2023; Karten et al., 2025; Zhang et al., 2024). Critics, by contrast, question whether such simulations are the “right thing” at all, warning against the risks of treating speculative outputs as predictive truths (Arnstein, 1969; Zhou et al., 2024; Agnew et al., 2024; Kapania et al., 2025; Li et al., 2025).
Our work suggests a third path: through iterative, stakeholder-engaged design, LLM agent simulations can become the right thing within particular policy contexts. Rather than assuming simulations are inherently useful or inherently useless, we found that usefulness emerges through a design process in which developers and policymakers jointly explore a design space defined by the intersection of technical possibilities and policy needs.
Indeed, by situating simulation development within the everyday decision-making practices of emergency preparedness professionals, we observed how LLM agent simulations could have practical utility. The simulations did not replace policymaker judgment or provide absolute predictive forecasts. Instead, they functioned as technology probes, surfacing tacit knowledge of the problem domain and enabling the testing of hypotheses that would be costly, or impossible to trial in the real world. Policymakers gradually progressed from validating simulations against mundane scenarios to using them for novel explorations, converting abstract outputs into concrete training protocols and evacuation procedures.
At the same time, caution is warranted. As trust in simulation outputs builds, there is a risk that model errors and biases risk will be overlooked. In particular, demographic biases in agent behavior present a two-sided challenge (Li et al., 2025). On the one hand, such biases can sometimes be intentionally leveraged to mirror real-world disparities and stress-test policies against them. On the other hand, uncritical reliance on these biased outputs—especially in policy decision-making—can reinforce inequities rather than mitigate them (Obermeyer et al., 2019). Developers and policymakers must therefore carefully distinguish between accepting bias to appropriately model reality and allowing bias to shape policy recommendations.
The significance of our work lies not in claiming that LLM agent simulations should universally guide policy implementation, nor that iterative design automatically guarantees usefulness. Instead, our study demonstrates how LLM agent simulations can become useful under specific design conditions. The central question is not whether such simulations are useful for policy, but how we can design them to be useful for policy. Our case demonstrates one such pathway: not inevitable or universally generalizable, but achievable through iterative, context-sensitive design.
7.2. How Can We Design LLM Agent Simulation That Are Useful for Policy?
While our findings build on standard HCI practice, that iterative stakeholder engagement maximizes our chances of making useful technology, this section highlights distinctive design conditions specific to LLM agent simulations for policy use. These conditions double as design implications, outlining how such simulations can be developed and deployed in ways that are genuinely useful for policy by supporting implementation. They are intended to guide developers, designers, and policymakers working at the intersection of HCI, AI, and public decision-making.

### v2

@@ 5.6. Fifth Iteration: Exploring Better Policy Implementation
In the final iteration, we used the validated simulation to explore concrete policy implementation options raised by policymakers. We examined operationally plausible variations in emergency type (severe weather and bomb threat), announcement specificity, and coordination strategies, including repositioning coordinators and opening an additional exit. Interventions were evaluated relative to baseline practices based on “evacuation efficiency”, measured as the time required for 80% of agents to evacuate.
This iteration, conducted in August 2025, produced three actionable proposals: (1) opening a previously unused exit at the northwestern corner of the stadium, (2) training coordinators to guide both nearby and distant groups during localized threats, and (3) differentiating evacuation procedures by emergency type and incorporating these distinctions into training materials.
Figure 5. Cumulative evacuation progress across policy alternatives in a severe weather scenario. The figure shows the simulated proportion of agents evacuated over time after an official severe weather announcement instructing evacuation from the stadium. Each curve represents the cumulative percentage of agents who have exited under the current procedure and three alternative strategies explored with the developed stadium system (Fig. 2). Vertical dashed lines indicate the time required to evacuate 80% of agents—a threshold that policymakers routinely use to assess evacuation effectiveness. Opening the northwest exit (green) significantly reduces evacuation time relative to the current process (yellow), corresponding to a 23.0% reduction at the 80% threshold, while repositioning coordinators or emphasizing the south exit yields only marginal gains. These comparative trajectories supported policymakers’ reasoning about which procedural changes were likely to meaningfully improve evacuation efficiency.
Line chart showing cumulative evacuation over time for four evacuation processes. The x-axis shows time in minutes and the y-axis shows the estimated proportion of occupants evacuated from 0% to 100%. Four colored curves represent the current process and three alternative processes: opening a northwest exit, emphasizing a south exit, and changing coordinator positions. All curves rise over time toward full evacuation. Vertical dashed lines mark the time at which 80% of occupants have evacuated for each process. The northwest exit condition reaches 80% evacuation earliest, approximately 23% faster than the current process, while the other alternatives show similar or slightly slower times than the current process.
Figure 5 illustrates how simulation outputs informed discussion around Proposal 1. Across scenarios, variations in coordinators’ positions — which was one of the primary interventions policymakers’ initially considered — did not substantially affect evacuation time. In contrast, opening the northwest exit significantly improved evacuation efficiency (23.0% relative to baseline), whereas other exits produced negligible gains due to local constraints. Policymakers interpreted the effectiveness of the northwest exit in relation to its proximity to a major congestion area between students and family members. This intervention was robust in both severe weather and bomb thread scenarios, except when the threat itself was located near that exit.
Following this iteration, the policymakers adopted Proposals 2 and 3, integrating revised coordinator guidance and differentiated emergency protocols into standard operating procedures. They noted that in man-made threat scenarios such as bomb or active-shooter incidents, operational authority shifts to local police, limiting their ability to fully implement some process changes. Proposal 1 entered a formal feasibility review, consistent with institutional procedures for infrastructure modifications. These outcomes were documented in the August 2025 after-action report, which committed to continued use of the simulation for preparedness planning and training.
6. Findings
We identified five mechanisms through which LLM agent simulations became useful for policy practice in our setting:
@@ 6. Findings
    The Validation Filter: Simulations are most useful in scenarios where simulated behaviors can be validated
@@
    The Trust Bootstrap: Trust from validating mundane scenarios extends to novel ones
@@
    The “Fix-It” Response: “Wrong” simulations surface tacit domain knowledge from policymakers
@@
Across these five mechanisms, policymakers were not primarily evaluating whether the simulator was “accurate” in predicting exact motion trajectories. Instead, they asked a more operational question: can this simulation be used to justify a concrete change in training, staffing, messaging, or infrastructure? In practice, simulation “usefulness” emerged when: it created a shared reference point that policymakers could check against something they already knew; it could be interrogated when it looked wrong; and, it could be iterated upon until it supported decisions that policymakers were actually empowered to make. These findings help explain why the same stochastic and generative affordances that make LLM simulations attractive for exploring “unknown unknowns” also make them difficult to adopt without an initial path to validation and trust calibration. While our findings are grounded in our engagement with just one institution, these mechanisms can reflect recurring conditions of policy-facing simulation work (See Section 7.1). We discuss each in turn below.
@@ 6.1. Validation Filter: Usefulness Requires Scenarios That Can be Validated
In our study, simulations were considered useful only when outputs could be validated against policymakers’ experience or ground-truth data—an implicit requirement that became visible through iteration. We call this the “validation filter”: usefulness depends on domains where well-understood scenarios can be jointly checked before extending to hypothetical emergencies. In this context, validation did not mean comparing simulation outputs to a single ground-truth value. Rather, policymakers needed to be able to check simulation behavior against something they already knew from experience, observation, or institutional records. This surfaced immediately in Iteration 1 (Sec. 5.2), where P3 doubted validation was feasible:
“It’s very hard for us to access social media data in previous emergencies, and some important communication may happen at private Discord channels… And then you have people in other countries who may be having a conversation that’s blowing up way beyond what is actually happening here… we don’t know, I mean, I don’t know how to communicate directly with some of our audiences overseas.”
@@
Commencement was therefore appealing not only because it was high-stakes, but because it recurred and could be directly observed. This made it possible to establish a shared sense of what “normal” looked like before using the simulation to explore more uncertain scenarios.
@@ 6.2. Trust Bootstrap: Trust from Validating Mundane Scenarios Extends to Novel Ones
The validation filter creates a paradox: simulations are most valuable for hard-to-observe events (Park et al., 2022, 2023; Binz et al., 2025), yet they are trusted only when grounded in well-understood scenarios. We found that this paradox could be addressed through a “trust bootstrap” process. In practice, this bootstrapping took the form of a gradual expansion in scope. Policymakers first accepted the simulator’s representation of routine movement patterns and spatial constraints. Once that foundation was established, they became willing to explore scenarios that layered additional uncertainty on top of this baseline, such as alternative exit strategies or hypothetical disruptions.
In our study, once trust was established through successful validation in mundane scenarios (in our case, attendees exiting the stadium after commencement), that trust could be extended—or “bootstrapped”—to more novel and speculative scenarios where robust validation was harder or practically impossible. This process has become especially salient with recent advances in LLMs, which allow simulation systems to adapt to new, natural-language scenarios while building on validated setups. This dynamic was reflected in P1’s comment during the fourth iteration (Sec. 5.5):
“…when you showed a baseline evacuation without anything major, it mirrored the 20 minutes of what you all experienced with us with just the regular evacuation. And your simulation mirrored what we experience every time we have it. So that gives us some baseline trust, if you will, to see that it is connecting to what experiences we already had.”
@@
We also found that the iterative process was essential because it created opportunities for policymakers to respond to simulations with “fix-it” reactions. Even when policymakers were deeply familiar with their domain, they did not initially know which details were critical for interpreting a simulation until they saw something that looked off. Incomplete, simplistic, or otherwise “wrong” simulations therefore served as productive technology probes (Hutchinson et al., 2003). By reacting to outputs that seemed unrealistic or missing, policymakers revealed hidden assumptions and unspoken expertise, which in turn guided refinements that brought the simulations closer to institutional relevance.
@@
For example, our original commencement evacuation simulation did not consider accessibility needs (Sec. 5.3). When observing that LLM agents move at the same speed, P5 immediately flagged this omission:
@@
“Did you think about like people with disabilities and they might want to exit like through the one that has like the ramp or anything like that? Did you guys think about those factors?”
@@
This requirement—accounting for accessibility situations—did not emerge until after policymakers saw a simulation that failed to include it.
Similarly, while we had incorporated friendship groups based on earlier discussions, the simulation initially modeled only students’ evacuation. After seeing this version, P4 immediately surfaced a critical missing dimension:
“You start thinking about the fact that in a graduation ceremony, for every student, there is potentially minimum of two family members that will be in attendance. And if you’re thinking through the concept of a friend group having a very strong connection, you’ve also got to put into consideration the fact that there is a family connection and they are not with their family in that graduation component.”
P5 built on this observation, noting that families would coordinate around shared landmarks:
“Like they’ll meet up at landmarks that the family knows… So like if your family’s not familiar with CMU and you tell them like let’s meet by the goal post that’s something visually that they can see.”
Despite being core to actual evacuation behavior, these family dynamics and landmark-based coordination strategies were never mentioned in initial requirements discussions. These insights rarely surfaced in interviews and instead emerged when policymakers confronted unrealistic outputs.
LLM agent simulations consist of two core elements: the agents themselves and the environments in which they interact (Sumers et al., 2023). In our early iterations, we focused primarily on designing agents and leveraging LLM capabilities. However, policymakers consistently pointed out that without sufficient nuance in the interaction environment, the simulations lacked credibility. When the interaction environment was under-specified, agents could still produce fluent and socially plausible behavior. However, this surface-level plausibility became a liability rather than an asset: because the interaction environment was under-specified, realistic-looking behavior gave the impression of policy relevance without clearly signaling its limits, making it difficult for policymakers to judge when simulation outputs could or could not inform operational decisions.
“We were thinking that this evacuation is due to severe weather, which would be a fairly benign reason. But it could also be because of a bomb threat, and that opens up a whole different conversation.”
Similarly, viewing the simulation prompted policymakers to recall physical details they had initially overlooked. As the simulation became more refined, P1 noted a missing element:
“One of the things that we’re not having in the model is the bleachers which is the section in the back, which has a lot more flexibility because we have less volunteers but several thousand people.”
In academic demonstrations, developers often concentrate on designing LLM agents. Our findings show that environment design is as consequential as agent design for policy use. Context reinforced utility by anchoring simulations in recognizable realities, and it sparked imagination about new applications. Missing or misaligned contexts could cause temporary dips in confidence, but these gaps also triggered “fix-it” responses that revealed additional requirements (see Sec. 6.3).
Our iterative process ultimately revealed a co-evolution between policy requirements and simulation capabilities. While we initially aimed to design simulations that could directly inform “policy,” policymakers clarified that our assumption was misplaced. P1 stated:
“These are process changes, not policy or concept of policy is different. Our policy is to implement safe procedures to make sure that we protect life, property. So, policy has a context of what we intend to do… So, our policy is the same. What we’re trying to do is to find ways to fulfill the policy, which is protecting life.”
This distinction matters because policy commitments remain stable while implementation processes adapt to context. It also clarified an important boundary on when simulations became actionable. Policymakers found the simulator most useful when it addressed decisions that fell within their direct authority, such as training practices, coordinator placement, or internal procedures. In contrast, simulation outputs were harder to act on when potential changes depended on external command structures, legal authority, or approvals beyond the institution. In these cases, the simulator still supported discussion and sensemaking, but it did not translate as directly into changes in policy implementation. More broadly, policymakers treated the simulator as a mutual learning device — useful once baseline trust existed, but always contingent on what institutional knowledge they could supply. As P1 later reflected:
@@ 6.4. The Details Matter: Contextual Nuance in Interaction Environments Enables Policy Use
Usefulness was therefore recursive: simulations supported policy implementation by helping policymakers refine operational procedures that realize stable policy goals, and those refinements in turn shaped what the simulator needed to represent next.
@@ 7.1. Are LLM Agent Simulations Useful for Policy?
Scholars remain divided on the usefulness of LLM agent simulations for policy. This debate echoes Bill Buxton’s distinction between “building the thing right” versus “building the right thing (Buxton, 2010).” While prior work has emphasized increasingly sophisticated agent architectures (Park et al., 2024; Sumers et al., 2023; Tang et al., 2025; Gao et al., 2023), critics warn that speculative simulations risk being mistaken for predictive truth (Arnstein, 1969; Zhou et al., 2024; Kapania et al., 2025).
Our engagement suggests that this debate cannot be resolved by evaluating simulation systems in isolation. Policymakers expressed dissatisfaction with existing tools — such as tabletop exercises and traditional agent-based models — that struggle to support reasoning about dynamic, interdependent action. For example, P1 explained:
“Typically, we’ll sit in a meeting with police, look at a map, and say things like, ‘I’ll send three officers here.’ But we don’t really go deeper into questions like: How long will that take? What are other agents doing at the same time? How do people react dynamically?”
Policymakers further noted that these tools often abstract away social meaning central to real-world decision-making, resulting in what P4 described as a “blank human” model:
“When we normally do tabletop exercises, we’re dealing with a kind of blank human. We don’t know who the person is, and everyone implicitly behaves the same way. But that’s not how real life works.”
The same limitations also apply to traditional ABMs. These models can represent evacuation geometry and physical flow more realistically than naive tabletop exercises (Joo et al., 2013; Rozo et al., 2019; Chen and Zhan, 2008; Helbing and Molnar, 1995), but they still fall short of capturing social and semantic factors central to real-world decision-making, such as communication, social ties, differentiated roles, and heterogeneous interpretations of procedures (Sun et al., 2016; DeAngelis and Diaz, 2019; Kostakos et al., 2010; Aguirre et al., 2011). These gaps motivate the exploration of alternative simulation paradigms, not as replacements for existing policy tools, but as complements that foreground communication, interpretation, and institutional roles.
Against this backdrop of institutional dissatisfaction with existing tools, LLM agent simulations offer a different design space: one that integrates language, social meaning, and role-specific reasoning into physically grounded scenarios. Our goal is not to claim that LLM agents surpass carefully tuned ABMs in navigation accuracy. Instead, the persistent challenges that limit the institutional adoption of existing simulation tools motivate exploration of alternative paradigms. However, LLM agent simulations are not inherently useful; their outputs are stochastic, semantically rich, and difficult to validate, making trust calibration a central design challenge. We found that LLM agent simulations can become useful for policy implementation when embedded in specific design conditions. Rather than emerging from predictive accuracy alone, usefulness developed through iterative engagement in which simulations were validated against mundane scenarios, used to surface tacit institutional knowledge, and gradually repurposed for training and operational planning. In this role, simulations functioned less as forecasting tools and more as technology probes that supported shared reasoning about complex scenarios.
Nonetheless, caution is warranted. As trust builds, model errors and demographic biases may be overlooked (Li et al., 2025). While such biases can be deliberately leveraged to stress-test policies against real-world behavioral disparities, uncritical reliance on biased outputs risks reinforcing inequities (Obermeyer et al., 2019), requiring a careful distinction between modeling bias and allowing it to shape policy decisions.
These observations also suggest that the usefulness of LLM agent simulations is not uniform across policy domains. Our findings help identify a set of enabling conditions under which such simulations are most likely to be worth the effort: (1) a recurring baseline scenario that can be observed directly; (2) decision levers that policymakers can realistically change; and (3) outcomes that can be discussed in operational terms rather than abstract social dynamics.
Policy domains that meet these characteristics may similarly support the gradual development of trust in LLM agent simulations. Beyond emergency preparedness, examples include managing passenger flow and staffing during recurring transit disruptions, coordinating staffing and bed allocation during predictable hospital demand cycles, or planning operations for large, regularly scheduled public events. In these settings, policymakers can observe baseline conditions, adjust concrete levers such as staffing, routing, or communication workflows, and meaningfully relate simulation outputs to plausible operational changes.
By contrast, domains such as long-term climate adaptation, national security strategy, or online misinformation governance often lack stable, observable baselines and place key decision levers outside the control of any single policymaking unit. In such contexts, simulations may still offer exploratory value, but translating simulation outputs into actionable institutional change may require forms of validation, interpretation, and organizational alignment beyond those examined in this work.
Rather than arguing that LLM agent simulations should universally guide policy, this work shows how usefulness can emerge through context-sensitive design aligned with institutional practices and constraints. The central question is therefore not whether LLM agent simulations are useful for policy, but how they can be designed to become useful for policy practice.
7.2. How Can LLM Agent Simulations Be Designed to Become Useful for Policy Practice?
Building on these boundary conditions, we articulate design process implications specific to LLM agent simulations in policy contexts. While our findings align with established HCI insights that iterative stakeholder engagement increases the likelihood of producing useful technologies, this section highlights three implications that are distinctive to LLM agent simulations when they are intended to support policy implementation rather than exploratory analysis alone. Grounded in the five insights identified in Section 6, these implications outline how such simulations can be developed and deployed in ways that are genuinely useful within real policy practices.